OpenStudy (haaileyyy):

Please Help I'm about to start a test over finding the greatest common factor of monomials and i was wandering if anyone could help me

OpenStudy (haaileyyy):

Find the greatest common factor of the following monomials : \[48b ^{2}\] \[21b\]

ganeshie8 (ganeshie8):

Is this question from your test ?

OpenStudy (haaileyyy):

yes

ganeshie8 (ganeshie8):

Isn't it against your school policy to take help from others during a test ?

OpenStudy (haaileyyy):

No, its credit recovery. Since it is all online and we don't have a teacher to help us or teach us they encourage us to find all the help we can.

OpenStudy (photon336):

so no one teaches it to you at all? wow...

ganeshie8 (ganeshie8):

Fair enough, I can teach you these, one moment please

OpenStudy (haaileyyy):

There is a lesson we can read through but it doesnt make sense to me. I am really bad at algebra

ganeshie8 (ganeshie8):

\[\color{blue}{48}b ^{2}\] \[\color{blue}{21}b\]

ganeshie8 (ganeshie8):

Step1 : Deal with the friendly numbers 48 and 21

ganeshie8 (ganeshie8):

Can you factor 48 ?

OpenStudy (haaileyyy):

1, 2, 3, 4, 6, 8, 12, 16, 24, 48 is that what you mean?

ganeshie8 (ganeshie8):

"factor" is just a fancy word it simply asks you to write 48 as 2*24

ganeshie8 (ganeshie8):

Notice that 2*24 equals 48

ganeshie8 (ganeshie8):

But why stop there when we can factor 24 further and write it as 2*12

ganeshie8 (ganeshie8):

48 = 2*24 = 2*2*12

ganeshie8 (ganeshie8):

can we factor that 12 again ?

OpenStudy (haaileyyy):

\[2\times6\] ?

ganeshie8 (ganeshie8):

Awesome! who said you're bad at arithmetic/algebra ? 48 = 2*24 = 2*2*12 = 2*2*2*6

ganeshie8 (ganeshie8):

you can guess what we're gona do next why leave 6 like that, factor it too

OpenStudy (haaileyyy):

\[2\times3\]

ganeshie8 (ganeshie8):

48 = 2*24 = 2*2*12 = 2*2*2*6 = 2*2*2*2*3

ganeshie8 (ganeshie8):

see if you're happy with below : \[\color{blue}{48}b ^{2} = 2*2*2*2*3*b*b\]

ganeshie8 (ganeshie8):

I've just replaced 48 by our factored form and \(b^2\) by \(b*b\) easy peasy so far

OpenStudy (haaileyyy):

I understand so far

OpenStudy (haaileyyy):

@ganeshie8

ganeshie8 (ganeshie8):

Excellent! let's do the exact same thing to the other monomial \(\color{blue}{21}b\)

ganeshie8 (ganeshie8):

Can you factor 21 ?

OpenStudy (haaileyyy):

I'm not sure if this is right but i know it equals 21. . \[3\times7\]

ganeshie8 (ganeshie8):

Perfect! \[\color{blue}{21}b = \color{blue}{3*7*b}\]

ganeshie8 (ganeshie8):

Now comes the tricky part

ganeshie8 (ganeshie8):

We're ready to eyeball the greatest common factor

OpenStudy (haaileyyy):

3?

OpenStudy (haaileyyy):

3b*

ganeshie8 (ganeshie8):

\[\color{blue}{48}b ^{2} = \color{blue}{2*2*2*2*3}*b*b\] \[\color{blue}{21}b = \color{blue}{3*7}*b\]

ganeshie8 (ganeshie8):

That's it! very good job in figuring that out! \(\color{blue}{3}\) is common from numbers \(b\) is common from variables so \(\color{blue}{3}b\) is the greatest common factor here

OpenStudy (haaileyyy):

Thank you so much

ganeshie8 (ganeshie8):

yw

OpenStudy (kittiwitti1):

\(\text{If you're done with the question, don't forget to close it}\) ☺

OpenStudy (kittiwitti1):

& good luck on your credit recovery! ♣